# Rotatable and Movable Antenna-Enabled Near-Field Integrated Sensing and Communication

### v3

@@ V-A Simulation Setup
We consider a global Cartesian coordinate system with the origin located at the BS. The center of TP and RP are set as $\mathcal{C}^{\text{t}}=[0,0,5+D/2]^{\mathsf{T}}$ and $\mathcal{C}^{\text{r}}=[0,0,5-D/2]^{\mathsf{T}}$, respectively. The MAs at the BS are confined to a square moving region, which is modeled as a rectangle within the planar local Cartesian coordinate system relative to the center of TP/RP, with local coordinates ranging from $[-D/2,D/2]\times[-D/2,D/2]$. The communication users are assumed to be randomly distributed in the $x-y$ plane, specifically $d_{k}\sim\mathcal{U}(15,25)$, $\theta_{k}=\frac{\pi}{2}$, and $\phi_{k}\sim\mathcal{U}(-\frac{\pi}{2},\frac{\pi}{2})$ for $k=[1,K]$. The sensing target’s position is set as $d_{0}=10$, $\theta_{0}=\frac{\pi}{3}$, and $\phi_{0}=\frac{\pi}{4}$. The simulation parameters are detailed in Table I, unless stated otherwise. It should be noted that due to the large aperture size, the target and the users are located within the Rayleigh distance.
@@
$A$ Size of MA elements $\frac{\lambda^{2}}{4\pi}$
@@
$\sigma_{k}^{2},\sigma_{0}^{2}$ Noise power of users& target -110dBm
@@
$a_{1}$ Individual learning factor in Algorithm 1 and Algorithm 2 1.4
$a_{2}$ Global learning factor in Algorithm 1 and Algorithm 2 1.4
$\omega_{\min}$ Minimum inertial weight in Algorithm 1 and Algorithm 2 0.4
$\omega_{\max}$ Maximum inertial weight in Algorithm 1 and Algorithm 2 0.9
@@ V-C Performance Comparison for Different Setups
Fig. 4 shows the communication performance comparison for different $C_{\max}$ with $\Gamma_{\min}=6$dB. As shown, the downlink sum-rate increases as the increase of $C_{\max}$ and gradually slows down, which can be regarded as a manifestation of the CRB-rate Pareto boundary [37]. In Fig. 4, the performance gain from rotations is more significant compared to the sensing-centric design because the effective aperture loss is mitigated by the rotations of TP and RP. The downlink sum-rate of the setup with transmit & receive RMA is substantially higher than transmit FPA & receive FPA. At $C_{\max}=1.4\times 10^{-6}$, the former achieves $78.7$% performance improvement over the latter. This indicates that the proposed BS design and algorithm are desirable and appealing when the ISAC network is heavily loaded.
@@ V-D Performance for Discrete Rotations
In practical deployment, continuous rotation of RMAs is difficult to realize since RMAs need to be mechanically moved by physical devices, such as a stepper motor that can only adjust the rotations of each RMA in discrete steps, which are usually specified by the motor used [15]. As such, we employ discrete rotations for the transmit RMA & receive FPA setup in Fig. 5 and Fig. 6. Specifically, we uniformly discretize the rotation angles with $2$ to $8$ bits by quantizing the optimal rotation angles obtained by Algorithm 1 and Algorithm 2. From Fig. 5(a)-5(c), it is obvious that quantization reduces the sensing performance. As the number of quantization bits decreases, the system performance exhibits a gradual degradation. With increasing $\Gamma_{\min}$, the sensing performance for target’s position information consistently declines, indicated by sustained increase of RCRB. This characterizes the Pareto boundary of sensing and communication performance under discrete transmit antenna rotation angle setup. As shown in Fig. 6, quantization has a more significant impact on the reduction of communication performance. As the number of quantization bits decreases, the downlink sum-rate also gradually decreases, even falling below the setup with transmit 1D continuous rotation RMA & receive FPA. Similarly, as $C_{\max}$ increases, the growth of the downlink sum-rate gradually slows down, which reflects the CRB-rate Pareto boundary under discrete transmit antenna rotation angle setup. Fig. 5 and Fig. 6 indicate that sufficient quantization level of RMA discrete rotation should be chosen based on the performance requirement and hardware conditions in both sensing-centric and communication-centric systems.
@@ VI Conclusion
In this paper, we investigated the RMA-enabled near-field ISAC system. We derived an accurate near-field channel model by introducing the effective aperture loss to the conventional spherical wave model and formulated two optimization problems for the sensing-centric and communication-centric designs, respectively. To solve the non-convex sensing-centric problem, we proposed an AO-based algorithm with the SDR method for optimizing the transmit beamforming and the PSO method for optimizing RMAs’ rotations and positions. Moving on to the communication-centric problem, we first rewrote it into a more tractable form and then proposed a similar algorithm to optimize the transmit beamforming and RMAs’ rotations and positions. Numerical results verified the effectiveness of the proposed algorithms and the advantages of the considered RMA-enabled ISAC network compared to conventional FPA-based systems. The proposed RMA-enabled near-field ISAC system is expected to perform a more accurate sensing and higher downlink communication rate by optimally designing the transmit beamforming and the RMAS’ positions and rotations.

### v4

@@ V-A Simulation Setup
We consider a global Cartesian coordinate system with the origin located at the BS. The center of TP and RP are set as $\mathcal{C}^{\text{t}}=[0,0,5+D/2]^{\mathsf{T}}$ and $\mathcal{C}^{\text{r}}=[0,0,5-D/2]^{\mathsf{T}}$, respectively. The MAs at the BS are confined to a square moving region, which is modeled as a rectangle within the planar local Cartesian coordinate system relative to the center of TP/RP, with local coordinates ranging from $D/2,D/2]\times[-D/2,D/2]$. The communication users are assumed to be randomly distributed in the xyx-y plane, specifically $d_{k}\sim\mathcal{U}(15,25)$, $\theta_{k}=\frac{\pi}{2}$, and $\phi_{k}\sim\mathcal{U}(-\frac{\pi}{2},\frac{\pi}{2})$ for $k=[1,K]$. The sensing target’s position is set as $d_{0}=10$, $\theta_{0}=\frac{\pi}{3}$, and $\phi_{0}=\frac{\pi}{4}$. The simulation parameters are detailed in Table I, unless stated otherwise. It should be noted that due to the large aperture size, the target and the users are located within the Rayleigh distance.
@@
AA Size of MA elements $\frac{\lambda^{2}}{4\pi}$[10]
@@
$\sigma_{k}^{2},\sigma_{0}^{2}$ Noise power of users& receiver -110dBm
@@
$a_{1}$ Individual learning factor in Algorithm 1 and Algorithm 2 1.4[40]
$a_{2}$ Global learning factor in Algorithm 1 and Algorithm 2 1.4[40]
$\omega_{\min}$ Minimum inertial weight in Algorithm 1 and Algorithm 2 0.4[40]
$\omega_{\max}$ Maximum inertial weight in Algorithm 1 and Algorithm 2 0.9[40]
@@ V-C Performance Comparison for Different Setups
Fig. 4 shows the communication performance comparison for different $C_{\max}$ with $\Gamma_{\min}=6$dB. As shown, the downlink sum-rate increases as the increase of $C_{\max}$ and gradually slows down, which can be regarded as a manifestation of the CRB-rate Pareto boundary [41]. In Fig. 4, the performance gain from rotations is more significant compared to the sensing-centric design because the effective aperture loss is mitigated by the rotations of TP and RP. The downlink sum-rate of the setup with transmit & receive RMA is substantially higher than transmit FPA & receive FPA. At $C_{\max}=1.4\times 10^{-6}$, the former achieves 78.778.7% performance improvement over the latter. This indicates that the proposed BS design and algorithm are desirable and appealing when the ISAC network is heavily loaded.
@@ V-D Performance for Discrete Rotations
In practical deployment, continuous rotation of RMAs is difficult to realize since RMAs need to be mechanically moved by physical devices, such as a stepper motor that can only adjust the rotations of each RMA in discrete steps, which are usually specified by the motor used [18]. As such, we employ discrete rotations for the transmit RMA & receive FPA setup in Fig. 5 and Fig. 6. Specifically, we uniformly discretize the rotation angles with $2$ to $8$ bits by quantizing the corresponding parts of each particle obtained by (38) in Algorithm 1 and Algorithm 2. From Fig. 5(a)-5(c), it is obvious that quantization reduces the sensing performance. As the number of quantization bits decreases, the system performance exhibits a gradual degradation. With increasing $\Gamma_{\min}$, the sensing performance for target’s position information consistently declines, indicated by sustained increase of RCRB. This characterizes the Pareto boundary of sensing and communication performance under discrete transmit antenna rotation angle setup. As shown in Fig. 6, quantization has a more significant impact on the reduction of communication performance. As the number of quantization bits decreases, the downlink sum-rate also gradually decreases, even falling below the setup with transmit 1D continuous rotation RMA & receive FPA. Similarly, as $C_{\max}$ increases, the growth of the downlink sum-rate gradually slows down, which reflects the CRB-rate Pareto boundary under discrete transmit antenna rotation angle setup. Fig. 5 and Fig. 6 indicate that sufficient quantization level of RMA discrete rotation should be chosen based on the performance requirement and hardware conditions in both sensing-centric and communication-centric systems.
@@ VI Conclusion
In this paper, we investigated the RMA-enabled near-field ISAC system. We derived an accurate near-field channel model by introducing the effective aperture loss to the conventional spherical wave model and formulated two optimization problems for the sensing-centric and communication-centric designs, respectively. To solve the non-convex sensing-centric problem, we proposed an AO-based algorithm with the SDR method for optimizing the transmit beamforming and the PSO method for optimizing RMAs’ rotations and positions. Moving on to the communication-centric problem, we first rewrote it into a more tractable form and then proposed a similar algorithm to optimize the transmit beamforming and RMAs’ rotations and positions. Numerical results verified the effectiveness of the proposed algorithms and the advantages of the considered RMA-enabled ISAC network compared to conventional FPA-based systems. The proposed system is expected to achieve more accurate sensing and higher downlink communication rate across a variety of scenarios. In particular, we foresee that the integration of RMA will open up new opportunities in massive MIMO, unmanned aerial vehicle communication, and adaptive multiple access. These directions deserve to be further explored in the future.